OpenStudy (anonymous):

convert to 51 celcius to degrees fahrenheit if If necessary, round your answer to the nearest tenth of a degree. c=5/9 (f-32) f=9/5 c+32

OpenStudy (jdoe0001):

seems to me you're set, so, do so then :)

OpenStudy (anonymous):

lol what? i need help i dont know what to do after

OpenStudy (jdoe0001):

you've got the formulas already, is all you really need

OpenStudy (anonymous):

no but i dont know what to do with them i dont know how to solve the problem im stuck on it

OpenStudy (jdoe0001):

celsius=5/9 (farenheit-32) farenheit=9/5 celsius+32

OpenStudy (jdoe0001):

so... still confused?

OpenStudy (anonymous):

i need to solve it put 51 c to fahrenheit but i dont know how to convert it

OpenStudy (anonymous):

thanks!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

just plug c=51 in f=(9/5)*c+32

OpenStudy (anonymous):

got it thank you so much!

OpenStudy (anonymous):

was confused for a while